how would i graph y-2=-1/4(x-2)?
do i have to move it to slope intercept form first?

Answer by Alan3354(69443)

how would i graph y-2=-1/4(x-2)?
It's not required, but it might make it easier to find points.
y-2=-1/4(x-2)?
y = (-1/4)*(x-2) + 2
It's a straight line, so 2 points are needed
Pick values for x, find y
x....y
0....2.5
2....2
Plot the points, draw a line thru them.